# BLACKOUT: Data-Oblivious Computation with Blinded Capabilities

### v3

@@ 9. Related Work
Today, the development of software that handles secret data, such as cryptographic libraries or operating-system kernels, must take precautions in order to avoid inadvertently introducing side-channel leakage. Most code is not naturally side-channel resistant as writing constant-time code is often counterintuitive for programmers who are accustomed to optimizing for performance or hardware resource cost (Intel, 2019). Similarly, common toolchains used for development are not well suited for constant-time programming for the following reasons: (1) optimizing compilers can break data-obliviousness properties of high-level code (Schneider et al., 2024), and (2) data-oblivious libraries, such as (Authors, 2025; agora lovecruft et al., 2025; McLean, 2023b; klutzy, 2015) must balance security with the cost disabling optimizations, e.g., by implementing large parts of constant-time functionality in architecture-dependent inline assembly blocks. A proper, complete solution to this problem would likely involve deep changes to LLVM compiler, like adding a field to LLVM’s type system to indicate which values require timing protection and severely restrict the program transformation performed on such values by optimization passes at a significant cost to performance (McLean, 2023b). Another approach used by constant-time Rust libraries takes advantage of the that inline assembly blocks are opaque to LLVM’s optimization passes (klutzy, 2015; McLean, 2023a). By implementing all constant-time operations in inline assembly, the compiler can be prevented from performing any optimizations whatsoever on constant-time code. However, such code cannot be analyzed for memory-safety correctness by the Rust compiler either, and inhibits even simple optimizations such as constant folding and algebraic simplification.
Researchers have been interested in the problem of writing and compiling data-oblivious and constant-time code (Molnar et al., 2005; Bernstein, 2005, 2006; Coppens et al., 2009; Cleemput et al., 2012; Andrysco et al., 2015; Rane et al., 2015; Ohrimenko et al., 2016; Fisch et al., 2017; Shaon et al., 2017; Zheng et al., 2017; Ahmad et al., 2018; Mishra et al., 2018; Sasy et al., 2018; Eskandarian and Zaharia, 2019; Borrello et al., 2021; ElAtali et al., 2024a). As a result, a variety of different checking tools that provide informal (Reparaz et al., 2017; Weiser et al., 2018; Wichelmann et al., 2018, 2022; Disselkoen, 2024; Langley, 2025) and formal (Cauligi et al., 2017; Daniel et al., 2020) guarantees have been proposed. An actively maintained online list of “constant-timeness’ verification tools (CT-tools) contains, at the time of writing, 55 different tools (Centre for Research on Cryptography and Security, 2025). However, testing whether a program is data-oblivious remains challenging as these tools are typically not integrated into modern toolchains, have significant technical limitations including high overheads to compilation time, and high-numbers of false-positives, and are difficult to use (Jancar et al., 2022; Geimer et al., 2023; Fourné et al., 2024). The principal shortcoming of static analysis approaches is that data-obliviousness can only be defined at machine-code level, rather than for high-level language constructs. The pessimistic view on the problem of achieving constant-time cryptographic implementations is that ensuring constant-time behavior through coding practices and alone does not work in general (Pornin, 2025), because optimizations employed by compilers continously increase in aggressiveness, and new compilers emerge in unexpected contexts, such as in-silicon just-in-time (JIT) compilers deep in CPU hardware. Moreover, a pervasive lack of documentation (Pornin, 2018b) throughout the whole software and hardware stack makes constant-time guarantees unachievable outside situations where the target hardware is extremely narrowly defined. Consequently, effectively enforcing data-obliviousness guarantees must be achieved through a hardware/software codesign (Liu et al., 2015; Yu et al., 2019; ElAtali et al., 2024b, a) which facilitates explicit contracts between programmers who have the domain-knowledge to annotate secret data their programs are intended to operate on, and the hardware which can apply that information to protect the secret data from leakage at the hardware level. Data-oblivious ISA (Yu et al., 2019; ElAtali et al., 2024b), discussed in § 2.3, operate under the underlying principle that hardware-based taint tracking is able to make side-channel leakage through timing from a silent failure of a systems security properties, into an explicit failure. Yet, programming for data-oblivious ISAs remains hard due to lack of explicit software/hardware contracts integrated into the toolchain.
@@
BLACKOUT addresses these shortcomings by introducing a hardware/software co-design that integrates seamlessly into existing mainstream toolchains. Unlike traditional static analysis tools,it provides architectural support through hardware-enforced taint tracking, which reliably converts implicit timing side-channel leaks into explicit faults. By augmenting the CHERI capability model with blinded capabilities, BLACKOUT ensures memory-safety and data-obliviousness without invasive architectural changes. It simplifies secure programming by modifying the CHERI Clang/LLVM compiler to infer when data-oblivious invariants should be enforced. Thus, unlike previous solutions which tackle these challenges one-by-one, BLACKOUT effectively reduces compiler-induced vulnerabilities and alleviates the burden on developers by addressing both usability and security challenges inherent in previous solutions.
@@ Acknowledgments
[…]

### v4

@@ 9. Related Work
Software that handles sensitive data—such as cryptographic libraries or operating-system kernels—must carefully avoid introducing side-channel leakage. Most code is not naturally resistant to such leaks as writing constant-time code is often counterintuitive for programmers used to optimizing for performance or resource efficiency (Intel, 2019). Development toolchains are not well suited for constant-time programming for two main reasons: (1) optimizing compilers can break data-obliviousness properties of high-level code (Schneider et al., 2024), and (2) constant-time libraries (Authors, 2025; agora lovecruft et al., 2025; McLean, 2023; klutzy, 2015) must balance security with the cost of disabling optimizations, e.g., by implementing constant-time functionality in architecture-specific inline assembly.
Verifying constant-time code.
Researchers have explored various methods for writing and verifying data-oblivious and constant-time code (Molnar et al., 2005; Bernstein, 2005, 2006; Coppens et al., 2009; Cleemput et al., 2012; Andrysco et al., 2015; Rane et al., 2015; Ohrimenko et al., 2016; Fisch et al., 2017; Shaon et al., 2017; Zheng et al., 2017; Ahmad et al., 2018; Mishra et al., 2018; Sasy et al., 2018; Eskandarian and Zaharia, 2019; Borrello et al., 2021; ElAtali et al., 2024a), resulting in numerous tools offering informal (Reparaz et al., 2017; Weiser et al., 2018; Wichelmann et al., 2018, 2022; Disselkoen, 2024; Langley, 2025) and formal (Cauligi et al., 2017; Daniel et al., 2020) guarantees. An actively maintained list of “constant-timeness” verification tools (CT-tools) currently includes 55 such tools (Masaryk University Centre for Research on Cryptography and Security, 2025). However, both static analysis or formal methods face signficant challenges. The principal shortcoming of static analysis approaches is that data-obliviousness can only be defined at machine-code level, rather than for high-level language constructs. Capturing microarchitectural subtleties of real-world hardware in formal models (and keeping the models up-to-date as hardware evolves) is difficult. Static analyses may not be sound and can lead to over- or under-tainting. Lastly, testing whether a program is data-oblivious remains challenging as tools built on static and formal analysis are typically not integrated into modern toolchains, have significant technical limitations including high overheads to compilation time, many false-positives, and are difficult to use (Jancar et al., 2022; Geimer et al., 2023; Fourné et al., 2024).
In practice, constant-time coding practices alone are unreliable (Pornin, 2025). Compiler optimizations continue to evolve, and new compilers emerge in unexpected contexts (e.g., in-silicon just-in-time compilers in CPU hardware). Documentation gaps across the software and hardware stack further hinder efforts, especially when target platforms are not narrowly confined (Pornin, 2018b). Thus, robust enforcement of data-obliviousness requires hardware/software codesign (Liu et al., 2015; Yu et al., 2019a; ElAtali et al., 2024b, a). Such designs allow programmers to annotate sensitive data, enabling hardware to enforce protection against side-channel leakage. Data-oblivious ISAs (Yu et al., 2019a; ElAtali et al., 2024b), discussed in § 2.3, use hardware-based taint tracking to transform silent timing side-channel leaks into explicit hardware faults. Yet, programming for data-oblivious ISAs remains hard due to the lack of integrated toolchain support for expressing such software/hardware contracts.
Architectural-Mimicry (AMi) (Winderix et al., 2024a) introduces new ISA primitives for more efficient control-flow-linearization, addressing how to make programs data-oblivious. AMi is complementary to data-oblivious ISAs, which enforce data-obliviousness through ISA contracts. However, it requires greater developer involvement as it requires developer write assembly manually (with correct usage of new primitives); incorrect usage can silently leak secrets.
@@
BLACKOUT complements static-analysis and formal verification approaches by bridging the gap between high-level verified data-oblivious software and the low-level machine code. BLACKOUT enforces data-oblivious implementation through a hardware/software co-design that integrates with mainstream toolchains. Unlike static analysis tools, it provides hardware-enforced taint tracking, reliably converting implicit timing leaks into explicit faults. By extending the CHERI capability model with blinded capabilities, BLACKOUT ensures both memory-safety and data-obliviousness without invasive architectural changes. It simplifies secure programming by modifying the CHERI Clang/LLVM compiler to infer when data-oblivious invariants should be enforced. Thus, BLACKOUT simultaneously addresses usability, security, and performance challenges that previous approaches tackled in isolation.
Securing speculation.
Many hardware defenses adapt CPU microarchitectures to defend against Spectre and other speculative execution side-channel attacks by isolating the microarchitectural elements, such as cache hierarchies  (Kiriansky et al., 2018; Yan et al., 2018; Sakalis et al., 2019; Saileshwar and Qureshi, 2019; Taram et al., 2019; Ainsworth and Jones, 2020; Kim et al., 2020; Zhao et al., 2020; Loughlin et al., 2021), that can either influence speculation or leak data across security domains. Speculative taint-tracking techniques(Yu et al., 2019b; Fustos et al., 2019; Schwarz et al., 2020; Daniel et al., 2023) delay instructions dependent on speculatively loaded secret data. Of recent work, ProSpeCT (Daniel et al., 2023) formalizes the constant-time policy with respect to control flow and memory accesses for a broad class of speculation mechanisms. Its scope is guaranteeing that hardware does not leak secrets of constant-time programs during speculation. Unlike data-oblivious ISAs, neither ProSpeCT nor other transient execution defenses enforce their guarantees on non-speculative execution. As such, they do not help developers write constant-time code. BLACKOUT prevents secret leakage even when programs are not constant-time and helps developers refine annotations. Future work can integrate formal approaches with BLACKOUT so that once developers transform code, it can be guaranteed data-oblivious.
@@ Acknowledgments
We thank our colleagues at University of Waterloo and Ericsson: Adam Caulfield, Christoph Baumann, Håkan Englund, Santeri Paa-
volainen, and Sini Ruohomaa for their comments on drafts of this paper. We also thank the University of Cambridge Computer Laboratory members, particularly Jessica Clarke, Jonathan Woodruff, Peter Rugg, Robert Watson, and former member Alex Richardson.
This work is supported in part by Natural Sciences and Engineering Research Council of Canada (grant number RGPIN-2020-04744), and the Government of Ontario (RE011-038). Views expressed in the paper are those of the authors and do not necessarily reflect the position of the funders.